### v1

# LazyEviction: Lagged KV Eviction with Attention Pattern Observation for Efficient Long Reasoning

Abstract
Large Language Models (LLMs) exhibit enhanced reasoning capabilities by employing Chain-of-Thought (CoT). However, the extended reasoning sequences introduce significant GPU memory overhead due to increased key-value (KV) cache size, particularly in tasks requiring long reasoning sequences, such as mathematics and programming. Existing KV cache compression methods mitigate memory bottlenecks but struggle in long reasoning tasks. In this paper, we analyze attention patterns in reasoning tasks and reveal a Token Importance Recurrence phenomenon: a large proportion of tokens receive renewed attention after multiple decoding steps, which is failed to capture by existing works and may lead to unpredictable eviction on such periodically critical tokens. To address this, we propose LazyEviction, a lagged KV eviction framework designed to maintain reasoning performance while reducing KV memory. LazyEviction is an Observation Window-based Lagged Eviction Mechanism retaining latent recurring tokens by performing lagged evictions across decoding steps, which contains two key components: (1) Recurrence Interval Tracking for capturing temporal variations in token importance, and (2) an Maximum Recurrence Interval-Centric Eviction Policy that prioritizes eviction based on tokens’ recurrence patterns. Extensive experiments demonstrate that LazyEviction reduces KV cache size by 50% while maintaining comparable accuracy on mathematics reasoning datasets, outperforming state-of-the-art methods. Our findings highlight the importance of preserving recurring tokens, which are critical for maintaining knowledge continuity in multi-step reasoning tasks.
1 Introduction
Large Language Models (LLMs) have emerged with advanced capabilities in complex reasoning tasks by enabling Chain-of-Thought (CoT) to elicit step-by-step inference wei2022chain ; kojima2022large . Recent advancements, such as OpenAI’s o1 jaech2024openai and DeepSeekR1 guo2025deepseek , further demonstrate that scaling up CoT lengths from hundreds to thousands of reasoning steps could continuously improve LLM reasoning. However, the increased reasoning sequences introduce substantial memory and computational overhead. Due to the autoregressive nature of LLM decoding, longer CoT outputs lead to proportional increases in GPU memory for caching key-value(KV) states. These issues become particularly pronounced when reasoning sequences extend into thousands of reasoning tokens, resulting in significant GPU memory costs. For example, in mathematics cobbe2021training ; hendrycks2021measuring and programming tasks chen2021evaluating , the reasoning sequences can grow up to 16k tokens, resulting in more than 100GB KV cache with a batch size of 32, which exceeds the memory capacity of even high-end GPUs.
KV cache compression has emerged as a promising approach to alleviate the memory bottleneck caused by increasing KV cache sizes. Unlike traditional long-context input tasks that focus on the prefilling phase li2024snapkv ; cai2024pyramidkv ; feng2024ada ; fu2024not , long-reasoning tasks perform long-context generation that requires dynamic KV cache compression in multiple decoding steps. The critical challenge lies in compressing generated KV caches while minimizing performance degradation. Existing approaches mitigate this through selective eviction strategies: early works liu2023scissorhands ; xiao2023efficient preserve subsets of recent and initial KV pairs, while others oren2024transformers ; chen2024nacl ; he2025treekv ; zhang2023h2o ; adnan2024keyformer leverage attention scores to prioritize critical tokens and evict others, which can be further divided into two types: (a) Current Attention-based Eviction (TOVA oren2024transformers , NACL chen2024nacl , TreeKV he2025treekv ) and (b) Cumulative Attention based Eviction (H2O zhang2023h2o , Keyformer adnan2024keyformer , MorphKV ghadia2025dialogue , RaaS hu2025efficient ). However, as shown in Fig. 1, these works are greedy eviction strategies at each decoding step, which discards tokens deemed temporarily unimportant while overlooking their potential future importance. We empirically observe that recurrence of token importance is a frequent phenomenon in multi-step reasoning processes, which typically involve conditional or summary information within reasoning chains, aligning with the capabilities in verification, backtracking, and summarization in reasoning models gandhi2025cognitive . The previous works approaches risk prematurely discarding crucial tokens during reasoning steps, ultimately sacrificing task accuracy.
To achieve efficient KV cache compression while preserving reasoning performance, we analyze attention patterns during decoding phases. Our investigation reveals a Token Importance Recurrence phenomenon: a large proportion of tokens in reasoning tasks receive renewed attention after multiple decoding steps. In this paper, we refer to such tokens as recurring tokens, where the attention score of recurring tokens may be low within an interval while suddenly increasing in the future. Empirical analysis demonstrates that these tokens often contain information vital for later reasoning stages, and premature eviction of recurring tokens leads to catastrophic performance degradation. Crucially, existing KV compression methods usually fail to identify such tokens due to inherent limitations.
Figure 1: Comparison of different KV Eviction methods. The dark squares represent that the token has a relatively higher attention score. (a) Current Attention-based Eviction execute stepwise evictions using immediate attention scores. (b) Cumulative Attention-based Eviction integrate historical attention for eviction decisions. Both (a) and (b) fails to preserve recurring tokens during their low-attention intervals. (c) LazyEviction performs lagged KV evictions based on the observation window to detect latent recurring tokens and prevent prematurely discarding them.
To address this challenge, we propose LazyEviction, a novel Observation Window-based Lagged Eviction Mechanism. Instead of per-step greedy decisions, we perform lagged KV evictions at every $W$ decoding steps to detect latent recurring tokens and prevent prematurely discarding them. There are two key componebts: (1) Recurrence Interval Tracking: For each token, we record its Maximum Recurrence Interval (MRI)—the longest observed period between its attention spikes during decoding to capture the temporal variation of token importance. (2) MRI-Centric Eviction Policy: When reaching memory budgets, we prioritize evicting tokens whose $\Delta T$ exceeds their MRI, where $\Delta T$ is the time elapsed since the token’s last important recurrence. Tokens with $\Delta T<MRI$ are retained as they still have the potential to become important in the future. LazyEviction shifts from step-wised greedy eviction to window-wised predictive retention: by continuously tracking MRI and executing lagged eviction during observation windows, we can effectively detect latent recurring tokens and prevent prematurely discarding them. To sum up, our key contributions are:
  1. 1.
    We observe Token Importance Recurrence in the reasoning task, where specific tokens exhibit recurrent attention patterns (i.e., being re-activated after multiple decoding steps). These recurring tokens often represent foundational formulations or intermediate conclusions which are crucial for maintaining knowledge continuity in multi-step reasoning.
  2. 2.
    We propose the LazyEviction, a Lagged KV Eviction scheme integrating temporal observation window and recurrence interval tracking. Our framework enables prediction of future token importance, thereby preventing premature eviction of recurring tokens.
  3. 3.
    We have implemented LazyEviction on the latest large reasoning model (e.g. DeepSeek-R1-Distill-Qwen and DeepSeek-R1-Distill-Llama). It has been verified on mathematics reasoning datasets that our method can maintain comparable performance while reducing 50% KV budget, surpassing the SOTA KV cache compression methods.
2 Related Works
Various approaches have been developed to improve LLM’s efficiency in handling long reasoning tasks. These methods can be classified into three main categories: KV cache compression, long reasoning compression, and system-level optimizations.
KV cache compression: Considering the inherent sparsity of attention mechanisms, these works focus on retaining critical tokens’ KV pairs while evicting non-essential ones within constrained memory budgets. Recent works focus on compressing KV cache during the model decoding phase. We summarize the existing works and analyze their limitations as follows:
  1. 1.
    Static Positional Retention: Methods like Scissorhands liu2023scissorhands and StreamingLLM xiao2023efficient employ static retention of initial and recent tokens under fixed budgets. While sustaining text coherence, their rigid retention policies cannot detect latent recurring tokens.
  2. 2.
    Current Attention-based Eviction: TOVA oren2024transformers , NACL chen2024nacl , and TreeKV he2025treekv execute stepwise evictions using immediate attention scores. This myopic strategy fails to preserve recurring tokens during their low-attention intervals.
  3. 3.
    Cumulative Attention-based Eviction: H2O zhang2023h2o , Keyformer adnan2024keyformer , and MorphKV ghadia2025dialogue integrate historical attention for eviction decisions. RaaS hu2025efficient extends this through dynamic updated timestamp. These approaches still mistakenly regard latent recurring tokens as unimportant due to continuous low historical scores.
Long Reasoning Compression: Recent works have focused on directly compressing the output of reasoning models. For instance, TALE han2024token and SoT aytes2025sketch adopts prompts engineering to control the length of CoT outputs. TokenSkip xia2025tokenskip achieves output compression through fine-tuning on condensed CoT datasets. InftyThink yan2025inftythink and LightThinker zhang2025lightthinker equip models with summarization capabilities to reduce token consumption per reasoning step. Approaches like COCONUT hao2022training , CCoT cheng2024compressed , and SoftCoT xu2025softcot enable reasoning directly in compressed latent spaces. Distinctively, LazyEviction maintains the original output length of LLMs while enhancing memory and computational efficiency through optimized KV cache usage, demonstrating compatibility with these compression approaches.
System-level Optimizations: By optimizing GPU memory usage and batch processing, vLLM kwon2023efficient , FlashAttention dao2022flashattention , and FlashDecoding hong2023flashdecoding imporve both time and memory efficiency in decoding. Additionally, approaches like QUEST tang2024quest and OmniKV hao2025omnikv store all KV caches in CPU and dynamically schedule them during decoding phase. Since these methods retain all KV caches, they can handle recurring tokens proposed in this paper. However, they incur non-negligible I/O latency from CPU-GPU transfers and rely on precise critical cache selection. By offloading non-critical tokens to CPU, LazyEviction can be compatible with such techniques.
3 Observations and Motivations
Figure 2: (a) shows the performance degradation for SOTA Methods on reasoning tasks. With the same KV cache compression ratio (e.g., 50%), compared with traditional language modeling tasks dataset PG-19 rae2019compressive , the performance of both H2O and TOVA has decreased by more than 20% on GSM8K dataset. (b) is the visualization of the importance variation by selecting Top-50% important tokens. Tokens at the same position show different importance at different decoding steps.
Figure 3: Visualization of Token Importance Recurrence. We observe attention maps across different heads of DeepSeek-R1-Distill-Llama-8B. We find most attention maps (>80%) show Token Importance Recurrence. (a) and (b) show an attention map with recurring tokens and their corresponding positions. (c) statistically analyzed the MRI distribution in different models among different tasks.
The goal of KV cache eviction is to identify and evict KV pairs that are non-critical for future decoding steps. To preserve the correctness of subsequent decoding results, it is critical to retain tokens with potential future importance. In this section, we first analyze the previous works and observe attention patterns in reasoning tasks. We summarize our findings as follows:
Finding 1: Previous KV cache compression algorithm, while demonstrating high performance in general long-context generation scenarios, exhibit degraded performance when applied to long reasoning tasks.
We compare the performance of H2O zhang2023h2o and TOVA oren2024transformers on reasoning tasks versus general long-context generation tasks in Fig.2(a), H2O and TOVA report competitive performance on standard language modeling dataset, but exhibit significant performance degradation on GSM8K under the same KV cache compression ratio. To diagnose this failure, we track the positions of top-50% important tokens across decoding steps in Fig.2(b) and observe that tokens critical in later steps are often missing in earlier steps. Due to their low importance scores in intermediate steps, these tokens are prone to premature eviction in prior works.
Finding 2: Token Importance Recurrence A prevalent phenomenon in reasoning tasks, where tokens initially exhibit strong attention weights, receive low attention weights during later decoding steps, and subsequently regain attention after a certain decoding steps interval. Recurring tokens contain critical information for the reasoning and prematurely evicting them will cause knowledge in-continuity in multi-step reasoning tasks.
[…]
Finding 3 Most recurring tokens exhibit an MRI far smaller than the generated length and can be detected via an observation window, which motivates our design of the Observation Window-based Lagged Eviction Mechanism.
As shown in in Fig. 3(c), We statistically analyzed the MRI distribution in different models among different tasks. As the output length of the model increases, the MRI gradually becomes larger, but most tokens’ MRI remian relatively small. For example, the output length of the Qwen model on the MATH500 dataset can reach 8k, but we have statistically found that the MRI of 80% of the tokens is less than 175.Therefore, as long as the window size exceeds the MRI of most tokens, the majority of recurring tokens will be identified.
4 LazyEviction
In this section, we introduce our LazyEviction, which is a novel framework of Observation Window-based Lagged Eviction. We first model LazyEviction as a multi-step dynamic optimization problem in Appendix A.1. Unlike existing works that perform KV eviction decisions at every decoding step, our LazyEviction adopts a window size $W$ as the interval for KV eviction decisions. As illustrated in Fig. 4(a), during decision-making, LazyEviction always retains the most recent $W$ KVs to preserve local coherence. When the number of KVs exceeds the predefined budget $B(B\gg W)$, LazyEviction executes KV eviction. The overall procedure of our LazyEviction is summarized in Appendix. A.3.
To select $B-W$ KVs from past caches to retain, LazyEviction tracks importance variations (i.e., MRI) of each token at every decoding step (Recurrence Interval Tracking). For eviction decisions, importance scores, which are defined to indicate tokens’ potential future importance, are calculated based on MRI. Consequently, LazyEviction selectively retains only the tokens in the KV cache most likely to be critical in the future (MRI-Centric Eviction).
Figure 4: Overview of our proposed LazyEviction Framework, where the dark squares represent that the token has a relatively higher attention score. (a) LazyEviction performs eviction decision at intervals of $W$ steps. The workflow contains two key operations: (b) Dynamic MRI Tracking according to updated important timestamp, and (c) MRI-Centric Scoring during decision phases, where tokens predicted to be critical for future steps are retained. (d) The MRI-Centric Score fundamentally predicts future token importance by analyzing historical patterns of importance variation (i.e., MRI and the time elapsed since their latest timestamp).
Recurrence Interval Tracking
To identify importance variations for recurring tokens, we employ timestamps to track the time of being important. Consistent with the definition in RaaS hu2025efficient , whenever a token receives an attention score exceeding a threshold $\alpha$ (e.g., $\alpha=0.0001$), its latest timestamp is updated to the value of current decoding step $t$. Thus, during the whole decoding phase, LazyEviction dynamically maintains a timestamp vector $\mathbf{TS}_{t}$ for all retained tokens, which records only the most recent activation time of tokens (illustrated in the leftmost figure in Fig. 4(b)).
However, only relying on $\mathbf{TS}_{t}$ fails to capture the temporal patterns of Importance Recurrence. To address this, we introduce the MRI to record the longest interval between two consecutive activations in history. For newly generated tokens, MRI is initialized to 0. The MRI value is updated as follows:
$\mathbf{MRI}_{t}=\max\{\mathbf{MRI}_{t-1},\mathbf{TS}_{t}-\mathbf{TS}_{t-1}\}$ (1)
MRI-Centric Eviction
When $\left|{\mathcal{S}_{t}}\right|$ exceeds $B$, KV eviction are triggered at periodic intervals $t=kW$ ($k\in\mathbf{N}^{+}$). Leveraging the tracked historical importance dynamics via $\mathbf{MRI}_{t-1}$ and $\mathbf{TS}_{t}$. We aim to predict each token’s future importance using an MRI-Centric Importance Score. As illustrated in Fig 4(d) the score should be designed under the following two intuitive heuristics:
  1. 1.
    H1: Given a token $i$, if time elapsed since its last activation ($t-T{S_{t}}[i]$) is greater than its $MR{I_{t}}[i]$, the possibility to be important in the future will decrease. The larger the ratio between $t-T{S_{t}}[i]$ and $MR{I_{t}}[i]$, the less likely token $i$ is important.
  2. 2.
    H2: Tokens with smaller MRI values have higher possibility to be important in the future, as frequently recurring tokens are deemed more critical. Notably, if $MR{I_{t}}[i]=0$, the token has never been activated since its generation (MRI is initialized as 0), resulting in an importance score of zero.
Based on the above intuitive heuristics, the MRI-Centric Importance Score can be formalized as:
[…]
We define ${}\exp\{\cdot\}$ term as H1-Score, which reflects the likelihood of the token regaining importance within the next observation window. Define $1/{MR{I_{t}}[i]}$ as H2-Score to prioritizes tokens with smaller MRI values. For each eviction decision, we first compute $\mathbf{I}_{t}$ with Eq. 2. Then, we retain the most recent $W$ KV pairs and the $\rm{Top}(B-W)$ KV pairs with the highest scores.
Disscussion about the choice of $W$ and $\alpha$
In LazyEviction, the observation window with size $W$ serves to detect recurring tokens. However, since the most recent $W$-th tokens must be preserved, the setting of $W$ inherently introduces a trade-off between local and global information retention. An excessively large $W$ risks losing critical past tokens, while an overly small $W$ may fail to detect numerous recurring tokens. To resolve this dilemma, we employ the observation from Fig. 3(c), setting $W$ as the MRI threshold corresponding to the 80% tokens. This threshold can be statistically determined through offline analysis by randomly selecting 50 samples. Our experiments in Sec. 5.3 analyze how $W$ affects reasoning performance.
Since the selection of $\alpha$ influences the temporal distribution of token timestamps, both extremely large and small $\alpha$ hinder effective identification of tokens’ importance variation. As the methodology for determining $\alpha$ has been thoroughly discussed in RaaS hu2025efficient , we directly adopt their approach in LazyEviction. We analyze the effect of setting different $\alpha$ in Sec. 5.3.
5 Experimental Results
In this section, we select DeepSeek-R1-Distill-Llama-8B  [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B  [^2] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B as the base models for evaluating the efficiency of LazyEviction. The evaluations are carried on two widely used mathematical reasoning benchmarks: GSM8Kcobbe2021training and MATH-500hendrycks2021measuring . We compare LazyEviction with FullKV and three representative KV cache compression methods, including TOVA oren2024transformers , H2O zhang2023h2o , and RaaS hu2025efficient . More implementation details can be found in Appendix A.2.
5.1 Performance Comparison with Baselines
Table 1 presents a comprehensive performance comparison of our LazyEviction and baselines. The experimental results demonstrate that LazyEviction consistently outperforms other methods across two datasets and two models, achieving performance comparable to the baseline FullKV with minimal deviation. In DeepSeek-R1-Distill-Llama-8B, at a KV cache compression ratio of 50%, our method maintains performance comparable to FullKV, while other methods exhibit significant degradation. Notably, the RaaS algorithm shows an accuracy drop of nearly 50% on the GSM8K dataset. In DeepSeek-R1-Distill-Qwen-7B, at a KV cache compression ratio of 65%, our method achieves an accuracy reduction of less than 2.1% compared to FullKV, outperforming other three algorithms.
Table 1: Performance of our LazyEviction compared with three baselines on the GSM8K and MATH-500 dataset with DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B. The best results among all methods are in bolded.
Base Model DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Qwen-7B
Dataset GSM8K MATH-500 GSM8K MATH-500
FullKV 74.68 62.0 91.74 86.4
KV Cache Comression Ratio = 50% KV Cache Comression Ratio = 65%
H2O zhang2023h2o 62.24 (-12.44) 56.0 (-6.0) 89.38 (-2.36) 81.2 (-5.2)
TOVA oren2024transformers 66.26 (-8.42) 56.2 (-5.8) 80.14 (-11.6) 79.0 (-7.4)
RaaS hu2025efficient 39.88 (-34.8) 58.2 (-3.8) 86.35 (-5.39) 83.0 (-3.4)
LazyEviction 72.70 (-1.98) 60.4 (-1.6) 90.07 (-1.67) 84.6 (-1.8)
Figure 5: Trade-off between accuracy and cache budget of LazyEviction and other baselines on (a) GSM8K and (b) MATH-500 datasets using DeepSeek-R1-Distill-Llama-8B Model, and (c) GSM8K and (d) MATH-500 datasets using DeepSeek-R1-Distill-Qwen-7B Model.
Trade-off between Accuracy and Cache Budget
We evaluating the performance of LazyEviction in optimizing the trade-off between KV cache budget and accuracy compared to baselines. As shown in Fig.  5, our method maintains higher accuracy across various budgets for both datasets and models. When the cache budget is large, other methods perform slightly worse than our approach. However, with a smaller cache budget, other methods experience substantial accuracy degradation. Notably, the RaaS algorithm experiences a pronounced accuracy drop at low cache budgets, as it prioritizes retaining prompt tokens, which limits the retention of other essential tokens and leads to poor performance. By implementing lagged KV eviction and leveraging MRI to track recurring tokens, our approach consistently achieves optimal performance, even under stringent budget constraints.
5.2 Memory Efficiency of LazyEviction
Figure 6: KV cache memory usage of different algorithms, including (a) the KV cache memory usage with varying output length (0-8k tokens) on DeepSeek-R1-Distill-Qwen-7B, and average KV cache memory usage on (b) DeepSeek-R1-Distill-Llama-8B and (c) DeepSeek-R1-Distill-Qwen-7B for GSM8K datasets under different cache budget.
We evaluate the KV cache memory usage of different algorithms. First, we evaluate the KV cache memory usage on the MATH-500 dataset using DeepSeek-R1-Distill-Qwen-7B, analyzing its variation with output token length. As shown in Fig. 6, the memory consumption of FULLKV increases linearly with the number of tokens. In contrast, the memory usage of TOVA, H20, and RaaS algorithms remains constant once the token count exceeds the budget. LazyEviction, with the observation window mechanism, exhibits minor fluctuations in KV cache memory usage after surpassing the budget, but Fig. 6 indicates that its memory growth is minimal compared to the other three algorithms. To further assess LazyEviction’s memory efficiency, we conduct experiments by calculating the average KV cache memory usage on the GSM8K dataset. Fig. 6 illustrates the performance of DeepSeek-R1-Distill-Llama-8B, demonstrating that our approach reduces cache memory to 55.85% while maintaining an accuracy drop of less than 1%. Similarly, Fig.  6 shows the performance of DeepSeek-R1-Distill-Qwen-7B, where our method achieves a 50.8% reduction in cache memory with an accuracy decrease of less than 5%.
5.3 Ablation Study
Table 2: Baseline algorithms with observation window mechanism, evaluating on the GSM8K dataset using DeepSeek-R1-Distill-Llama-8B with $B=192$ and $W=25$.
Model DS-Distill-Llama-8B
Budget = 192
LazyEviction 73.62
H2O 66.26
+ window (W=25) 69.98 (+3.72)
TOVA 70.43
+ window (W=25) 71.34 (+0.91)
RaaS 54.36
+ window (W=25) 58.21 (+3.85)
Lagged Eviction with an Observation Window
Since our method retains a slightly increased number of KV pairs within the observation window, which may lead to the perceived performance gains being predominantly attributed to these additional tokens. To eliminate the impact of the additional tokens within observation window, we explore the effects of other algorithms with the observation window mechanism on performance. As shown in Table 2, when H20, TOVA, and RaaS algorithms perform KV eviction every $W$ decoding steps, their performance improves. This improvement arises from the linear increase in retained KV pairs between consecutive decision decoding steps. However, their performance remains inferior to that of LazyEviction. LazyEviction employs MRI to track recurring tokens, effectively capturing their potential future importance, which leads to superior performance.
The Selection of Window Size $W$.
We investigate the impact of the window size $W$ on DeepSeek-R1-Distill-Llama-8B performance across two datasets in Table 3. As $W$ increases, the model’s accuracy generally improves. This is because a larger window size allows the LazyEviction mechanism’s MRI metric to detect more recurring tokens, resulting in better capturing a token’s potential future significance, thereby enhancing performance. However, when the window size becomes too large, LazyEviction risks discarding critical past tokens, leading to a slight performance decline. Thus, the choice of $W$ presents a trade-off between retaining local and global information.
Table 3: Performance of different setting of the window size $W$ on DeepSeek-R1-Distill-Llama-8B.
Dataset GSM8K (Budget=192)
FullKV W=4 W=8 W=16 W=25 W=32
Accurracy 74.68 68.69 70.13 70.28 73.61 73.18
Dataset MATH-500 (Budget=256)
FullKV W=8 W=16 W=32 W=52 W=64
Accurracy 62.0 54.4 54.6 55.2 57.6 57.2
The Selection of $\alpha$
We further explore the impact of the threshold $\alpha$ on the GSM8K dataset using two models. The value of $\alpha$ influences the MRI of each token, with $\alpha=0.0005$ producing optimal results for DeepSeek-R1-Distill-Llama-8B and $\alpha=0.0001$ yielding the best performance for DeepSeek-R1-Distill-Qwen-7B, as detailed in Table 4. First, when $\alpha$ is small, tokens are more readily flagged as important, but the MRI of these tokens is reduced, which does not accurately reflect the periodicity of their significance. Second, when $\alpha$ is large, the steps where tokens gain importance are not captured, preventing the detection of their recurring patterns.
Table 4: Performance of different setting of $\alpha$.
Model DS-Distill-Llama-8B (Budget=192)
FullKV $\alpha$=0.0001 $\alpha$=0.0005 $\alpha$=0.001
Accurracy 74.68 72.48 73.61 73.23
Model DS-Distill-Qwen-7B (Budget=768)
FullKV $\alpha$=0.00001 $\alpha$=0.0001 $\alpha$=0.001
Accurracy 91.74 90.44 90.75 90.67
Table 5: Performance of using different importance score.
LazyEviction 73.62
w/o the H1-Score 69.67
w/o the H2-Score 73.23
The Discussion of the Importance Score
We further investigate the influence of the MRI-Centric Importance Score, as shown in Table 5. If the importance score $\mathbf{I}_{t}$ calculation excludes the H1-Score, accuracy significantly decreases, underscoring the effectiveness of the recurrence interval tracking mechanism. Similarly, omitting the H2-Score leads to a slight accuracy drop, indicating that tokens with smaller MRI values are indeed more likely to become important in the future. Collectively, these findings demonstrate the efficiency of LazyEviction’s MRI-Centric Eviction Policy.
6 Limitations
While LazyEviction significantly improves KV cache efficiency for long reasoning tasks, several limitations remain. Firstly, our evaluation primarily focuses on mathematical reasoning, futher evaluation is required on programming and other types of reasoning datasets. Second, fixed-length observation windows limits LazyEviction’s flexibility across tasks and models. Future work should investigate adaptive window size strategies suitable for diverse reasoning scenarios. Third, although empirically validated, the conditions under which Token Importance Recurrence occurs lack formal theoretical characterization, particularly in complex reasoning pathways.
7 Conclusion
In this work, we address the critical bottleneck of KV cache memory overhead in long reasoning tasks. Our analysis reveals the Token Importance Recurrence phenomenon, where certain tokens receive renewed attention after multiple decoding steps. Our experiment shows the significant of these recurring tokens in reasoning tasks. Based on this insight, we propose LazyEviction, a novel lagged eviction framework that introduces temporal observation windows, recurrence interval tracking, and an MRI-centric eviction policy to preserve recurring tokens while dynamically managing memory usage. Experimental results demonstrate that LazyEviction achieves a 50% reduction in KV cache size while maintaining comparable performance, outperforming SOTA methods. By shifting from greedy eviction to predictive retention, our work provides a new paradigm for efficient KV cache compression in long-context generation scenarios.
Appendix A Technical Appendices and Supplementary Material
A.1 Mathematical Formulation of LazyEviction
Suppose the sequence length is $N_{t}$ at decoding step $t$. We denote $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$ and $\mathbf{V}_{i}$ as the query, key and value vectors of the $i$-th token. The attention output at the decoding step $t$ can be formulated as:
$\mathbf{A}_{t}({\mathbf{Q}_{{N_{t}}}},\mathbf{K},\mathbf{V})={\rm{Softmax}}\left({\frac{{\mathbf{Q}_{N_{t}}{\mathbf{K}^{T}}}}{{\sqrt{{d_{h}}}}}}\right)\mathbf{V},$ (3)
where $\mathbf{K}=[\mathbf{K}_{1},...,\mathbf{K}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$, $\mathbf{V}=[\mathbf{V}_{1},...,\mathbf{V}_{N_{t}}]\in\mathbb{R}^{N_{t}\times d_{h}}$ and $d_{h}$ is the head dimension. We define $\mathbf{S}_{t}$ as the set of the KV pairs $(\mathbf{K}_{i},\mathbf{V}_{i})$ for all the previous tokens $i<N_{t}$ cached in GPU memory at decoding step $t$. Since the size of $\mathbf{K}$ and $\mathbf{V}$ grows linealy with the decoding step $t$, current methods try to evict the unimportant KV pairs from $\mathcal{S}_{t}$ at each step $t$. Unlike the previous KV compression methods, we make KV eviction decisions at decoding step $t=kW$ ($k\in\mathbf{N}^{+}$). Denote $\mathcal{S}^{*}_{t}$ as the optimal KV pairs selected from $\mathcal{S}_{t}$ under the cache budget $B$. Thus, the KV eviction problem can be formulate as:
${\mathcal{S}_{t}^{*}=\mathop{\arg\min}\limits_{\left|{\mathcal{S}_{t}^{\prime}}\right|=B,{\mathcal{S}_{t}^{\prime}}\subseteq{\mathcal{S}_{t}}}}&{{{\left\|{{\mathbf{A}_{t}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K},\mathbf{V})-{\mathbf{A}_{t}^{\prime}}({\mathbf{Q}_{{N_{t}}}},\mathbf{K^{\prime}},\mathbf{V^{\prime}},{\mathcal{S}^{\prime}_{t}})}\right\|}_{2}}}\end{array}$ (4)
where $\mathbf{K^{\prime}}$ and $\mathbf{V^{\prime}}$ are selected KV pairs given by $\mathcal{S}^{\prime}_{t}$, $\mathbf{A}_{t}^{\prime}$ is the attention output computed by Eq. 3 after KV eviction. To minimize the effect of the KV eviction process, our goal is to find the optimal $\mathcal{S}^{*}_{t}$ to minimize the change in attention output.
It is intractable to directly solve Eq. 4 because of its combinatorial nature. We design a heuristic method to efficiently find the suboptimal solution for $\mathcal{S}^{*}_{t}$. First, as emphasized in previous works xiao2023efficient ; zhang2023h2o ; ghadia2025dialogue , retaining recent tokens is critical to ensuring coherence in the generated context. And to observe the importance variations of the newly generated tokens, we consistently preserve the latest $W$ KV pairs. Second, among the remaining past KVs, we retain $B-W$ KV pairs by prioritizing tokens most likely to exhibit importance recurrence in future steps. This selection is guided by an auxiliary score vector $\mathbf{I}_{t}$, which quantifies the potential future importance of tokens. Formally, our eviction policy is defined as:
$\mathcal{S}_{t}^{\prime}=\rm{Top}_{[B-W]}(\mathbf{I}_{t})\cup\mathcal{W}_{t}$ (5)
where $\rm{Top}_{[k]}(\cdot)$ represents the selection of top-k entries (KV pairs) from $\mathcal{S}_{t}$ and $\mathcal{W}_{t}$ is the set of the most $W$ recent KV pairs in $\mathcal{S}_{t}$.
A.2 Experiment Setup
Models and Datasets
By fine-tuning with CoT data generated by DeepSeek-R1, even smaller distilled models can acquire long-reasoning capabilities similar to DeepSeek-R1. Thus, we select DeepSeek-R1-Distill-Llama-8B  [^3] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B  [^4] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B as the base models for evaluating KV cache compression. The evaluations are carried on two widely used mathematical reasoning benchmarks: GSM8Kcobbe2021training and MATH-500hendrycks2021measuring . GSM8K is a high-quality, linguistically diverse dataset of elementary-level math problems requiring multi-step reasoning and basic arithmetic operations to solve. MATH-500 contains 500 challenging problems drawn from high school math competitions, spanning domains such as algebra, geometry, and number theory.
Baselines
To validate the performance of LazyEviction, we compare it against Full Cache and three representative KV cache compression methods, including TOVA oren2024transformers , which retains tokens with the highest attention scores at each decoding step; H2O zhang2023h2o , which preserves heavy hitter tokens (identified by cumulative attention scores) and recent tokens during decoding; and RaaS hu2025efficient , which selects tokens with the newest timestamps at every decoding step.
Implementation Details
We implement LazyEviction in HuggingFace transformers library  wolf2020transformers to integrate it into the existing attention mechanism. We set the observation window size equal to the value of 80% MRI threshold based on the preliminary observations (Fig. 3(c)), and the number of recent tokens in H2O is equal to LazyEviction’s window size. Unless stated otherwise, we set $\alpha=0.0005$ for DeepSeek-R1-Distill-Llama-8B model and $\alpha=0.0001$ for DeepSeek-R1-Distill-Qwen-7B model. Experiments are conducted on NVIDIA V100 (32GB) GPUs. We use Ubuntu 20.04 with Linux kernel 6.8.0 and CUDA 12.8. For the budget setting in Table. 1, we first statistic the average length of different models on two datasets (shown in Table. 6). Then we adopt the compression ratio to compute the cache budget $B$.
Table 6: Token Length Statistics of Different Models
DeepSeek-R1-Distill-Llama-8B
GSM8K MATH-500
Prompt Output Length Total Length Prompt Output Length Total Length
85.97 257.79 343.76 95.82 583.7 679.52
DeepSeek-R1-Distill-Qwen-7B
Prompt Output Length Total Length Prompt Output Length Total Length
118.31 922.43 1040.74 128.26 2855.45 2983.71
A.3 Overall Procedure of LazyEviction
We provide the pseudo-code of LazyEviction in Algorithm. 1, which contains Recurrence Interval Tracking at each step $t$ and makeMRI-Centric Eviction when reach the decoding step $t=kW$ ($k\in\mathbf{N}^{+}$). Thus, Lazyeviction tasks the current decoding step $t$, observation window size $W$, cache budget $B$, $\alpha$, and $\mathbf{A}_{t}$, $\mathbf{K}$, $\mathbf{V}$ as input.
Algorithm 1 LazyEviction
0:  $\mathbf{A}_{t}$, $\mathbf{K}$, $\mathbf{V}$, $W$, $B$, $\alpha$, $t$.
1:  for  All token $i$ do
2:     if $A_{t}[i]>=\alpha$ then
3:        Update $TS_{t}[i]=t$.
4:     end if
5:  end for
6:  Update $\mathbf{MRI}_{t}$ according to Eq. 1.
7:  Calculate $\mathbf{I}_{t}$ according to Eq. 2.
8:  if $t==kW$ and $B>=\left|{\mathbf{K}}\right|$ then
9:     Update KV pairs according to Eq. 5.
10:  end if
11:  return  $\mathcal{S}_{t}^{\prime}$.